OpenStudy (anonymous):

If f(x)= 2x 2 + 3, then f(x - 1)= 2x 2 + 4x + 5. True or false

OpenStudy (anonymous):

assume the 2x 2 = 2x^2 ?

OpenStudy (anonymous):

so false?

OpenStudy (anonymous):

mlhuyk it would be helpful if you used 2x^2 for 'two x squared'

OpenStudy (anonymous):

it is false... f(x) = 2x^2 + 3... f(x-1) = 2(x-1)^2 +3 = 2(x^2 -2x +1) +3= 2x^2 -4x +5...